A chemist has a solution of 50% pure acid and another of 80% pure acid. How many gallons of each will make 600 gallons of 72% pure acid?
Answer by addingup(3677)

.5x + .8(600 - x) = .72(600)
.5x + 480 - .8x = 432
-.3x = -48
x = 160 gallon of the 50% solution must be added to 600-160= 440 gallon of the 80% solution